OpenStudy (anonymous):

i dont know how to find the derivative of absolutevalue(x-4)

OpenStudy (anonymous):

it is 1 if x is bigger than 4, -1 if x is less than 4, and it is not differentiable at x = 4

OpenStudy (anonymous):

that is because \[f(x) = |x-4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq 4 \\ - x - 4& \text{if} & x < 4 \end{array} \right. \]

OpenStudy (anonymous):

hah got to use "piecewise"

OpenStudy (anonymous):

showoff

myininaya (myininaya):

x-4 if x>4

OpenStudy (anonymous):

stolen from zarkon(tm)

myininaya (myininaya):

-(x-4)=-x+4 if x<4

OpenStudy (anonymous):

actually just googled and got it, but zarkon put in the text

OpenStudy (anonymous):

@myininaya i wrote it correctly. oh yes!

myininaya (myininaya):

\[f(x) = |x-4| = \left\{\begin{array}{rcc} x - 4 & \text{if} & x \geq 4 \\ - x + 4& \text{if} & x < 4 \end{array} \right.\]

myininaya (myininaya):

no! i win

myininaya (myininaya):

tell him zarkon tell him i win

OpenStudy (zarkon):

you always win

myininaya (myininaya):

:)

OpenStudy (anonymous):

good answer

OpenStudy (anonymous):

grrr

OpenStudy (anonymous):

oh right! i did make a mistake. this is my template \[f(x) = |x+4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq -4 \\ - x - 4& \text{if} & x < -4 \end{array} \right. \]

OpenStudy (anonymous):

i copy and paste this in, but i didn't make all the necessary changes

myininaya (myininaya):

sure...

OpenStudy (anonymous):

ask zarkon it is the one i stole from him

myininaya (myininaya):

whatever makes you feel better about the mistake

OpenStudy (anonymous):

(pout)

myininaya (myininaya):

i'm sure you will bust me soon lol

OpenStudy (zarkon):

\[f(x) = |x-4| = \left\{\begin{array}{rcc} x - 4 & \text{if} & x > 4 \\ 0 & \text{if} & x=4\\ - x + 4& \text{if} & x < 4 \end{array} \right.\]

myininaya (myininaya):

fancy

OpenStudy (zarkon):

yes...ty ;)

OpenStudy (zarkon):

\[f'(x) =\left\{\begin{array}{rcc} 1 & \text{if} & x > 4 \\ \text{DNE} & \text{if} & x=4\\ -1& \text{if} & x < 4 \end{array} \right.\]

OpenStudy (anonymous):

so which piecewise is it? theoen with equal to or the one with DNE?

myininaya (myininaya):

what?

myininaya (myininaya):

this last thing zarkon wrote is the answer

myininaya (myininaya):

the deivaitve of |x-4| does not exist at x=4 but exist everywhere else

OpenStudy (zarkon):

or the first thing satellite73 wrote

myininaya (myininaya):

for x>4 f'=1 for x<4 f'=-1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

where did the one come from?

OpenStudy (zarkon):

\[\frac{d}{dx}(x-4)=1\] \[\frac{d}{dx}(-x+4)=-1\]

OpenStudy (anonymous):

oh okay understood